OpenStudy (anonymous):

We can draw two tangnets From the point p to circle |z + 5 + 2i| =2 the coordinates of p?

OpenStudy (anonymous):

\[\sqrt{(a+5)^2+(b+2)^2}=2\] For the equation of the circle, with z=a+ib. p is just anywhere outside the circle.

OpenStudy (anonymous):

@hartnn @experimentX @eliassaab @Ishaan94

OpenStudy (anonymous):

Did nt Get u @henpen

OpenStudy (anonymous):